OpenStudy (anonymous):

Still need help with this problem I found the slope which is -1/3 but I do not know how to put it into slope intercept form perpendicular to the line -3x+y=-30 and contains the point 9,-4

OpenStudy (anonymous):

so now you knwo that you line passes through (9,-4) hence its eq can be given by (y - (-4)) = m (x-9) where m is the slope you have found

OpenStudy (anonymous):

i hope you can do the rest..

OpenStudy (anonymous):

slope intercept form of the line y=mx+c so y - (-4)) = m (x-9) can also be written as y = m (x-9) -4

OpenStudy (anonymous):

Thanks for your time and giving me steps, I have several other problems just like this one to finish.

OpenStudy (anonymous):

yw hope u will be able to solve and undersatnd them as well